# OpenGraph: Towards Open Graph Foundation Models

### v2

@@ Appendix A Appendix
A.1. Model Optimization
@@
(8) $\mathop{\arg\min}_{\mathbf{\Theta}_{f}}\sum_{\mathcal{G}_{s}}\sum_{\bar{\mathcal{E}}_{s}\in\mathcal{G}_{s}}\mathcal{L}\Big{(}f(\mathcal{G}_{s}-\bar{\mathcal{E}}_{s},\phi;\mathbf{\Theta}_{f}),\leavevmode\nobreak\ \bar{\mathcal{E}}_{s}\Big{)}+\lambda\|\mathbf{\Theta}_{f}\|_{\text{F}}^{2}$
@@
Here, $\mathcal{G}_{s}-\bar{\mathcal{E}}_{s}$ represents the input graph $\mathcal{G}_{s}$ with the label edge set $\bar{\mathcal{E}}$ removed. All training graphs $\mathcal{G}_{s}$ are jointly trained with random alternations. To enhance the model’s adaptability to different graph projections $\phi$, we regenerate the projection function $\phi$ for each training graph $\mathcal{G}_{s}$ every 10 training steps. The weight $\lambda$ is used for the parameter decay regularization term.
A.2. Experiments
Table 4. Statistics of experimental datasets.
Dataset # Nodes # Edges # Features # Classes
Link OGBL-ddi 4,267 1,334,889 0 N/A
OGBL-collab 235,868 1,285,465 128
ML-1M 9,746 720,152 0
ML-10M 80,555 7,200,040 0
Amazon-book 144,242 2,380,730 0
Node Cora 2,708 10,556 1433 7
Citeseer 3,327 9,104 3,703 6
Pubmed 19,717 88,648 500 3
Generated Gen0 46,861 454,276 0 N/A
Gen1 51,061 268,007 0
Gen2 32,739 240,500 0
A.2.1. Experimental Datasets
  • OGBL-ddi. This dataset is used for drug-drug interaction prediction. Each node represents an approved or experimental drug. The edges represent the combined effect of taking two drugs together, which differs significantly from taking them individually.
  • OGBL-collab. This is an academic social relation dataset. Its nodes represent scholars, and edges denote collaborations. Each node is combined with a 128-dimensional average word embedding calculated from the author’s historical publications.
  • Movielens-1M & Movielens-10M. These two datasets are collected from the movie rating platform Movielens. The graphs are constructed by connecting users with the movies they have rated. They contain 1 million and 10 million rating records, respectively.
A.2.2. Evaluation Protocol
This section includes detailed description for our cross-dataset zero-shot setting and the few-shot settings for baselines. It also introduces our evaluation metrics.
Zero-shot setting. For our OpenGraph, we utilize a zero-shot learning setting in which OpenGraph is not trained on any of these real-world datasets but is tested using the training set information as input information, including the graph structures, node features, and node labels in the training set. To effectively generalize to unseen node labels in node classification with zero shot, taking inspiration from previous works [27], we treat the label classes as new nodes and connect the vanilla nodes with training labels to the corresponding class nodes. This strategy removes the requirement for learning class-related parameters in the zero-shot learning setting. This enhancement is also applied to baselines methods.
Evaluation metrics. In link prediction, we follow existing works [33] to conduct the full-rank test for each node. To be specific, for each node, all nodes not connected to it in the training set are ranked by the model. The top-N nodes are taken as positive predictions, and we calculate Recall@N scores with $N=20,40$. In node classification, we employ the widely-used Accuracy and Macro-F1 metrics [4].
A.2.3. Implementation Details
The reported results are obtained by pretraining our OpenGraph using three generated datasets: Gen0, Gen1, and Gen2. The statistics of these datasets are presented in Table 4. We first generate the Gen0 dataset without injecting the graph’s topological pattern. Subsequently, we generate Gen1 and Gen2 based on Gen0 by incorporating the graph pattern. In comparison to Gen1, the Gen2 dataset undergoes an additional densification process, where nodes with less than 10 edges are removed. To acquire the nodes for the Gen0 dataset, we prompt the LLM to iterate through all products on an e-commerce platform, with a maximum generation depth of $5$. The Gibbs sampling algorithm is initialized with nodes having $6$ random edges. To ensure low overlap between consecutive samples, we introduce a separation of $1000$ sampling steps before generating each new sample. The dynamic probability normalization maintains the last $T^{\prime}=5000$ sampling instances. The node locality incorporation involves using $7$ locality indices and $0.95$ decay rate.
A.2.4. Baselines
We give detailed descriptions for the baseline models in this section. 9 models from 4 categories are utilized.
Graph-agnostic Approaches.
  • MF. This is the matrix factorization approach which learns node embeddings to reconstruct the observed adjacency matrix. For the node classification task, we adapt it to learn embedding vectors for each node with the goal of predicting node labels.
Non-pretraining Graph Neural Networks.
  • GCN [16]. This approach utilizes iterative graph convolutional operators to extract the high-order topological information.
  • GAT [30]. This graph attention network learns weights for node-wise connections using the attention mechanism, to facilitate adaptive graph information propagation and aggregation.
  • GIN [40]. This method enhances the representation power of GNNs by employing a distinct graph encoding method that emphasizes the discrimination of non-isomorphic structures.
Graph Pre-training Models.
  • GraphCL [52]. This baseline method utilizes pre-training of graph models through the application of a self-discriminative contrastive learning task on learned node embeddings. It incorporates various graph augmentation techniques such as node drop, edge permutation, random walk, and feature masking.
  • DGI [31]. This method introduces a self-supervised pre-training task that aims to maximize the mutual information between the local node view and the global graph view.
Graph Prompt Tuning Methods.
  • GraphPrompt [21]. This work presents a unified framework for pre-training and prompt tuning of graph models. It introduces a learnable prompt layer that automatically identifies crucial information in the pre-trained model for downstream tasks.
  • GPF [5]. This is a universal graph prompt tuning framework designed for various graph pre-training strategies. It introduces two versions of a learnable graph prompt layer.
A.2.5. Alternative Graph Projection Methods (RQ2)
  • Degree embeddings. Degree embeddings are a commonly used node representation strategy for non-attributed graphs. Each degree number is assigned an independent embedding vector, and each node is initially represented by its degree representation.
A.2.6. End-to-end Training Performance (RQ6)
Table 5. Performance comparison with models trained on few-shot datasets, in terms of Recall@20 (%).
Dataset GCN GAT GIN OpenGraph
1-shot 5-shot 1-shot 5-shot 1-shot 5-shot 1-shot 5-shot
OGBL-ddi 2.79 7.05 5.80 7.11 5.30 7.35 7.93 8.60
ML-1M 4.32 10.54 2.45 15.06 4.66 14.58 16.58 19.19
Amazon-book 0.96 2.51 0.47 2.28 0.69 2.52 2.96 3.10
Figure 4. LLM prompt template and generation examples.
A.3. Graph Generation Algorithm
A.3.1. Prompt Template and Generation Examples
@@
A.3.2. Node Generation Algorithm
@@ 10             $\hat{\mathcal{V}}+=$ DivideNode($v^{\prime}$, $n+1$)
Algorithm 1 Node generation based on tree-of-prompt.
A.3.3. Edge Generation Algorithm
@@ 19             $\mathcal{I}+=[\textbf{a}^{t}\oplus v_{t^{\prime}}]$
Algorithm 2 Edge generation based on Gibbs sampling.

### v4

@@ Appendix A Appendix
The appendix offers additional details about our OpenGraph framework. It covers the neural model component, the generation algorithm, as well as the experimental settings and supplementary results.
A.1 Methodology
A.1.1 Zero-Shot Node Classification
Node classification tasks face challenges when transferring trained classification capabilities from one graph to another due to the heterogeneity of node classes across datasets. To address this challenge, our OpenGraph transforms the graph-specific classification task into a unified link prediction task. This task involves predicting links between regular nodes and special nodes representing different classes. By leveraging the generalized topology extraction capability and learning from observed class-node relations, our OpenGraph enables zero-shot node classification.
A.1.2 Handling Node Features
In attributed graphs, node attributes can vary across different graphs, including textual, numerical, and categorical features. To address the semantic differences in node features, our OpenGraph transforms these attributes into a unified graph structure format. This format can be easily tokenized by our unified graph tokenizer and comprehended by our trained graph transformer. In our approach, we sample node pairs with the highest similarity scores as augmented edges. The similarity scores $s_{i,j}$ between a node pair $(v_{i},v_{j})$ are calculated as $s_{i,j}=\textbf{f}_{i}^{\top}\textbf{f}_{j}$. Using these similarity scores, we select the top $B\times K$ edges for each batch, where each batch contains $B\times|\mathcal{V}|$ candidate edges. This strategy is applied to all experimental datasets that have node features.
A.1.3 Details of Scalable Graph Transformer
In the efficient self-attention with anchors, we determine the number of anchors $S$ by $S=d/H<B$, to ensure that the self-attention module incurs similar memory costs as other fully-connected components. Here, $H$ represents the number of attention head. More specifically, the self-attention process for each head can be summarized as follows:
$\textbf{e}_{t}^{(3)}$ $=\sum_{v_{a}}\alpha_{t,a}\textbf{W}^{(v)}\textbf{e}_{a}^{(2)},\leavevmode\nobreak\ \textbf{e}_{a}^{(2)}=\sum_{v_{t}}\alpha_{a,t}\textbf{W}^{(v)}\textbf{e}_{t}^{(1)}$
$\alpha_{t,a}$ $=\text{softmax}\Big{(}\frac{(\textbf{W}^{(q)}\textbf{e}_{t})^{\top}\cdot(\textbf{W}^{(k)}\textbf{e}_{a})}{\sqrt{d/H}}\Big{)}$ (7)
Our efficient self-attention involves embeddings $\textbf{e}_{*}^{(1)},\textbf{e}_{*}^{(2)},\textbf{e}_{*}^{(3)}$ for anchor nodes $v_{a}$ and vanilla nodes $v_{t}$. After each attention calculation, the results from multiple heads are concatenated, passed through a learnable linear layer, and connected with a residual connection. The parameters $\textbf{W}^{(q)},\textbf{W}^{(k)},\textbf{W}^{(v)}$ are the parameters of the attention layer. To reduce computational complexity, we employ a two-stage self-attention process. It transforms the $3\times B$-length sequence to a shorter $S$-length sequence and then reverses the process.
After the self-attention module, each layer of our scalable graph transformer includes a two-layer fully-connected block with residual connections, accompanied by two layer normalization modules. To ensure numerical stability, per-layer scaling is applied by element-wisely dividing embeddings by a selected constant $K=10$.
A.1.4 Model Optimization
@@
$\mathop{\arg\min}_{\mathbf{\Theta}_{f}}\sum_{\mathcal{G}_{s}}\sum_{\bar{\mathcal{E}}_{s}\in\mathcal{G}_{s}}\mathcal{L}\Big{(}f(\mathcal{G}_{s}-\bar{\mathcal{E}}_{s},\phi;\mathbf{\Theta}_{f}),$
¯$\bar{\mathcal{E}}_{s}\Big{)}+\lambda\cdot\|\mathbf{\Theta}_{f}\|_{\text{F}}^{2}$ (8)
@@
Here, $\mathcal{G}_{s}-\bar{\mathcal{E}}_{s}$ represents the input graph $\mathcal{G}_{s}$ with the label edge set $\bar{\mathcal{E}}$ removed. All training graphs $\mathcal{G}_{s}$ are jointly trained with random alternations. To enhance the model’s adaptability to different graph projections $\phi$, we regenerate the projection function $\phi$ for each training graph $\mathcal{G}_{s}$ every 10 training steps. $\lambda$ denotes the weight for $L_{2}$ regularization.
A.2 Graph Generation Algorithm
A.2.1 Prompt Template and Examples of Generated Nodes
@@
Figure 4: Prompt template and generation examples.
A.2.2 Node Generation Algorithm
@@ 10             $\hat{\mathcal{V}}+=$ DivideNode($v^{\prime}$, $n+1$)
Algorithm 1 Tree-of-prompt algorithm for node generation.
A.2.3 Edge Generation Algorithm
@@
A.3 Experiments
Table 4: Statistics of experimental datasets.
Dataset # Node # Edge # Feat # Class
Link OGBL-ddi 4,267 1,334,889 0 N/A
OGBL-collab 235,868 1,285,465 128
ML-1M 9,746 720,152 0
ML-10M 80,555 7,200,040 0
Amazon-book 144,242 2,380,730 0
Node Cora 2,708 10,556 1433 7
Citeseer 3,327 9,104 3,703 6
Pubmed 19,717 88,648 500 3
Gen. Gen0 46,861 454,276 0 N/A
Gen1 51,061 268,007 0
Gen2 32,739 240,500 0
A.3.1 Experimental Datasets
  • OGBL-ddi. This dataset is used for drug-drug interaction prediction. Each node represents a drug. The edges represent the combined effect of taking two drugs together, which differs significantly from taking them individually.
  • OGBL-collab. It is an academic social relation dataset. Its nodes represent scholars, and edges denote collaborations. Each node is combined with a 128-dimensional average word embedding calculated from the author’s publications.
  • Movielens-1M & Movielens-10M. These two datasets are both collected from the movie rating platform Movielens. The graphs are constructed by connecting users with the movies they have rated. The two datasets contain 1 million and 10 million rating records, respectively.
A.3.2 Evaluation Protocol
This section includes detailed description for our cross-dataset zero-shot setting and the few-shot settings for baselines. It also introduces the evaluation metrics used in our experiments.
Zero-shot setting. For our OpenGraph, we utilize a zero-shot learning setting in which OpenGraph is not trained on any of these real-world datasets but is tested using the training set information as input information, including the graph structures, node features, and node labels in the training set. To effectively generalize to unseen node labels in node classification with zero shot, taking inspiration from previous works Sun et al. (2022), we treat the label classes as new nodes and connect the vanilla nodes with training labels to the corresponding class nodes. This strategy removes the requirement for learning class-related parameters in the zero-shot learning setting. This enhancement is also applied to baselines methods.
Evaluation metrics. In link prediction, we follow existing works Wei et al. (2024) to conduct the full-rank test for each node. To be specific, for each node, all nodes not connected to it in the training set are ranked by the model. The top-N nodes are taken as positive predictions, and we calculate Recall@N scores with $N=20,40$. In node classification, we employ the widely-used Accuracy and Macro-F1 metrics Chen et al. (2022b).
A.3.3 Implementation Details
The reported results are obtained by pretraining our OpenGraph network using three generated datasets: Gen0, Gen1, and Gen2. The statistics of these datasets are presented in Table 4. We first generate the Gen0 dataset without injecting the graph’s topological pattern. Subsequently, we generate Gen1 and Gen2 based on Gen0 by incorporating the graph pattern. In comparison to Gen1, the Gen2 dataset undergoes an additional densification process, where nodes with less than 10 edges are removed. To acquire the nodes for the Gen0 dataset, we prompt the LLM to iterate through all products on an e-commerce platform, with a maximum generation depth of $5$. The Gibbs sampling algorithm is initialized with nodes having $6$ random edges. To ensure low overlap between consecutive samples, we introduce a separation of $1000$ sampling steps before generating each new sample. The dynamic probability normalization maintains the last $T^{\prime}=5000$ sampling instances. The node locality incorporation involves using $7$ locality indices and $0.95$ decay rate.
A.3.4 Baselines
We give detailed descriptions for the baseline models in this section. 9 models from 4 different research lines are utilized in our evaluation.
Graph-agnostic Approaches.
  • MF. This is the matrix factorization approach which learns node embeddings to reconstruct the observed adjacency matrix. For the node classification task, we adapt it to learn embedding vectors for each node to predict node labels.
Non-pretraining Graph Neural Networks.
  • GCN Kipf and Welling (2017). This approach utilizes iterative graph convolutional operators to extract the high-order topological information.
  • GAT Veličković et al. (2018a). This graph attention network learns weights for node-wise connections using the attention mechanism, to facilitate adaptive graph information propagation.
  • GIN Xu et al. (2018). This method enhances the representation power of GNNs by employing a distinct graph encoding method that emphasizes the discrimination of non-isomorphic structures.
Graph Pre-training Models.
  • GraphCL Zhu et al. (2021). This baseline method utilizes pre-training of graph models through the application of a self-discriminative contrastive learning task on learned node embeddings. It incorporates various graph augmentation techniques such as node drop, edge permutation, random walk, and feature masking.
  • DGI Veličković et al. (2018b). This method introduces a self-supervised pre-training task that aims to maximize the mutual information between the local view and the global view.
Graph Prompt Tuning Methods.
  • GraphPrompt Liu et al. (2023). This work presents a unified framework for pre-training and prompt tuning of graph models. It introduces a learnable prompt layer that automatically identifies crucial information in the pre-trained model to facilitate downstream tasks.
  • GPF Fang et al. (2023). This is a universal graph prompt tuning framework designed for various graph pre-training strategies. It introduces two versions of a learnable graph prompt layer.
A.3.5 Alternative Projection Methods (RQ2)
  • Degree embeddings. Using degree embeddings for node representation is a commonly used strategy for non-attributed graphs. Each degree number is assigned an independent learnable embedding vector, and each node is initially represented by its degree representation.
A.3.6 End-to-end Training (RQ6)
Table 5: Performance comparison with models trained on few-shot datasets, in terms of Recall@20 (%).
Model GCN GAT GIN OpenGraph
shot 1 5 1 5 1 5 1 5
ddi 2.79 7.05 5.80 7.11 5.30 7.35 7.93 8.60
ML1M 4.32 10.54 2.45 15.06 4.66 14.58 16.58 19.19
Amazon 0.96 2.51 0.47 2.28 0.69 2.52 2.96 3.10
@@ 19             $\mathcal{I}+=[\textbf{a}^{t}\oplus v_{t^{\prime}}]$
Algorithm 2 Edge generation algorithm.